OpenStudy (anonymous):

5g nacl are dissolved in 1000g water. the mole fraction for nacl and water are 0.15and0.85 b,0.00154and0.9984 c,0.001and0.999 d,0.0015and0.9985

OpenStudy (anonymous):

First determine the moles of NaCl and H_2O. \[5.00gNaCl\frac{1molNaCl}{58.44gNaCl}=0.0855molNaCl\] \[1000gH_2O\frac{1molH_2O}{18.0152gH_2O}=55.50molH_2O\] Now we total the number of moles: \[55.50mol+0.0855mol=55.59mol\] Now we determine the fractional percentages: \[\frac{0.0855}{55.59}=0.00154\] \[\frac{55.50}{55.59}=0.9984\] So option b is what you are looking for. I'd like to note that this answer didn;t follow proper SFs. If it had none of the answers presented would have worked. I had to adjust the 5g to 5.00g.